# A Systematic Survey on Large Language Models for Algorithm Design

### v1

@@ 3.3. LLMs as Extractors (LLMaE)
LLMaE employs LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then strategically utilized in the enhancement of algorithm-based problem solving (Fig 7). This process leverages the advanced natural language understanding capabilities of LLMs, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional analytical methods (Wu et al., 2024c). By integrating these insights into algorithm design, developers can create more precise and efficient solutions tailored to specific needs.
@@
LLMs have demonstrated exceptional capabilities in feature extraction across various domains for algorithm design, surpassing traditional heuristics and conventional deep learning models. For instance, Kristiadi et al. (2024) explore the use of LLMs in accelerating principled Bayesian optimization in the molecular space. LLMs are used as fixed feature extractors for standard Bayesian optimization surrogate models, and parameter-efficient finetuning methods are leveraged to obtain the posterior of the LLM surrogate. The experiments show that LLMs can be useful for Bayesian optimization over molecules, but only if they have been trained or finetuned with domain-specific data.
Beyond feature extraction, LLMs excel in text comprehension and knowledge extraction, when compared to conventional network models. An example in LLM-assisted algorithm selection by Wu et al. (2024c) utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. Extensive experiments validate the performance superiority of the proposed model, showcasing the efficacy of each key module. In another innovative application, Du et al. (2024c) propose a novel approach using a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem and decreasing energy consumption and AI model implementation costs. Additionally, Becker et al. (2024) discuss the use of LLMs in the automotive and supplier industries, specifically focusing on retrieval-augmented generation systems to improve information retrieval from technical documentation. Memduhoğlu et al. (2024) use LLM embeddings to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
The multi-modal capabilities of LLMs also distinguish them from traditional feature extraction methods. Park et al. (2024) demonstrate that incorporating lexical information extracted from LLMs into an acoustic-based speaker diarization system through a probabilistic multi-modal decoding process and joint acoustic and lexical beam searches can significantly improve speech-processing performance.
@@ 3.4. LLMs as Designers (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 8). This utilization of LLMs extends their application beyond traditional boundaries, enabling them to actively participate in algorithm development by generating heuristics, writing code snippets, or formulating functions that can be integrated into larger systems. By doing so, LLMs can significantly accelerate the algorithm design process, reduce human effort, and bring creativity and optimization to algorithm development, which is difficult to achieve through conventional methods alone.
@@
Function design might be one of the most straightforward application domains. Eureka (Ma et al., 2024b) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for reinforcement learning. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. In tests across 29 different reinforcement learning environments, including 10 types of robots, Eureka surpassed human experts in 83% of the tasks, with an average improvement of 52%. Another study on automatic reward function design (Li et al., 2024b) leverages LLMs to automatically design dense reward functions for reinforcement learning agents in environments with sparse rewards, termed Auto MC-Reward. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. FunSearch (Romera-Paredes et al., 2024) adopts LLMs for function generation in an evolutionary framework with a multi-iland population management. FunSearch demonstrated promising results on both mathematical problems and optimization problems.
@@
AEL (Liu et al., 2023b) presents an early attempt to adopt LLM as a designer to design an algorithm idea and its code implementation. It adopts LLM in an evolutionary framework to create, combine, and revise the algorithmic heuristic for the traveling salesman problem. The results show a significant improvement over simple greedy algorithms and robustness across different sizes. EoH (Liu et al., 2024d) is the new version of AEL with five prompt strategies for heuristic space searching. EoH is applied to the key heuristic component design for guided local search. The results outperform human-crafted metaheuristics and deep-learning-based neural solvers. Similar frameworks have been applied to cost-aware acquisition function design for Bayesian optimization (Yao et al., 2024a), adversary attack algorithm design for image adversary attack (Guo et al., 2024c), and heuristic design for edge server task scheduling Yatong et al. (2024). LLaMEA van Stein and Bäck (2024) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (Ye et al., 2024a) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. With the reflection, better results on black-box cases (no problem-specific information is provided in the prompts) have been observed. Moreover, Yan et al. (2024) explore the use of LLMs to generate drafts of high-quality algorithms for electronic health records phenotyping, aiming to improve efficiency and scalability in observational health research and Hazra et al. (2024) combine LLMs with heuristic planning to generate cost-effective plans, showing superior performance in LLM planning approaches.
Unlike previous studies on LLMaD that focus on optimizing a single performance criterion, MEoH (Yao et al., 2024b) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space, which can hardly addressed by conventional MOEAs.
@@
LLM-based agent design has also gained much attention. For example, ADAS (Hu et al., 2024) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents, leveraging the Turing Completeness of programming languages to explore novel prompts, tool use, and control flows. They present a novel algorithm, Meta Agent Search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains, highlighting their robustness and potential for general application. Further studies on LLM-based agent systems are discussed in (Sun et al., 2024a) and (Liu et al., 2024e).
@@
There have been significant advancements in using LLMs for optimization modeling. Ahmed and Choudhury (2024) evaluate the performance of various LLMs in converting linguistic descriptions into mathematical optimization problems. The study introduces LM4OPT, a fine-tuning framework for smaller LLMs that reduces the performance gap between smaller and larger models, with GPT-4 achieving a high accuracy on the NL4Opt dataset without relying on additional named entity information. Similarly, Tang et al. (2024a) identify four critical requirements for the training dataset of operational research LLMs and develops a semi-automated process called OR-Instruct to create synthetic data tailored to these requirements. They introduce the IndustryOR benchmark, the first industrial benchmark for testing LLMs on real-world OR problems. By applying the data from OR-Instruct to various open-source LLMs, the authors demonstrate significantly improved capabilities for optimization modeling. Moreover, Mostajabdaveh et al. (2024) introduce a novel multi-agent modeling framework for translating natural language descriptions into optimization models. Instead of relying on the solver execution, it adopts relations identifier agents and a multi-agent verification mechanism for effective assessment.
@@ 4. Search Methods
In this section, we investigate two branches of the interaction of LLMs with search methods: 1) search methods are designed algorithms and 2) search methods are LLM-based searching frameworks for algorithm design. In the first category, search methods are the target of LLMs, examples including LLMs for evolutionary algorithms and Bayesian optimization. In the other category, LLMs are used in the search methods to be iteratively queried to design better algorithms. This section categorizes the works according to the type of search methods without distinguishing the two branches.
@@ 4.1. Sampling
The most straightforward search manner is by repeatedly instructing LLM to sample new designs (Zhou et al., 2023). The best sample is selected as the final design. However, recent studies have shown that simple sampling from LLMs is insufficient in algorithm design (Zhang et al., 2024b).
@@ 4.1.1. Beam Search
Beam search explores a limited set of promising paths, or “beams”, in the search space rather than only one. Beam search has been widely used in diverse domains. For example, LLM-based prompt engineering (Zhou et al., 2023; Pryzant et al., 2023; Ma et al., 2024d) leads to promising results, e.g., (Zhou et al., 2023) demonstrates cases where when the beam search size increases from 4 to 128, the performance beats the human expert.
Except for prompt engineering, Lin et al. (2023) introduce Text2Motion, a language-based planning framework that enables robots to execute complex sequential manipulation tasks. It adopts beam search over tokens and enhances task planning with LLMs using feasibility heuristics. In contrast, Hazra et al. (2024) adopt beam search over actions. It introduces ”SayCanPay” to integrate the world knowledge of LLMs with heuristic planning principles to produce grounded and efficient action sequences, demonstrating superior performance in extensive evaluations compared to other LLM planning methods. Moreover, Park et al. (2024) use beam search with lexical information from LLMs for enhancing speech-processing tasks and Wei et al. (2024) propose a confidence-aware sub-structure Beam Search as a decoding method operating at the sub-structure level for structured data generation.
@@ 4.1.2. MCTS
MCTS is a tree search algorithm that uses random sampling to build a search tree and make decisions based on the results of these samples. It is particularly effective in problems with large and complex search spaces. For example, Dainese et al. (2024) propose GIF-MCTS, a code generation strategy with Generate, Improve and Fix using MCTS, which outperforms baselines on the Code World Models Benchmark and other benchmarks. In GIF-MCTS, nodes in the tree are programs and edges are actions. Each action taken from a parent node produces a new complete program, which is split into a “state” part and a “rollout” part, and stored in a child node. The upper confidence bound for trees is used to select which action to take. Similarly, VerMCTS (Brandfonbrener et al., 2024) constructs a search tree with progressive widening to effectively manage large action spaces defined by lines of code. In this search tree, expansion and evaluation are guided by the verifier, serving as a computationally inexpensive upper bound on the value function in the code synthesis MDP. Moreover, Wang et al. (2023b) regard prompt optimization as a strategic planning challenge, utilizing a principled planning algorithm based on MCTS to effectively explore the expert-level prompt space. Drawing inspiration from human trial-and-error learning, the proposed PromptAgent generates precise insights and detailed instructions by analyzing model errors and providing constructive feedback.
@@ 4.2.1. Hillclimb
Hillclimb search iteratively searches for better results, where the new one is generated based on the last one and the better one is survived. A basic version of Hillclimb is investigated in paper by Zhang et al. (2024b) for algorithm design. It achieves competitive results when compared to population-based evolutionary algorithms. The reasoning over the existing status is usually adopted to enhance each search step. For example, Li et al. (2024b) leverage LLMs to automatically design dense reward functions for reinforcement learning agents in environments with sparse rewards, termed Auto MC-Reward, which iteratively refines the reward function based on feedback from the agent’s interactions with the environment and Yang et al. (2023) develop a multi-module framework with innovative feedback mechanisms, demonstrating through both LLM-based and expert evaluations that LLMs can effectively produce scientific hypotheses that are both novel and reflective of reality.
@@ 4.2.2. Neighborhood Search
Neighborhood search methods are optimization techniques that explore the solution space by iteratively modifying a current solution to find an improved one in a structured neighborhood. Notably, the LLM-GS framework (Liu et al., 2024f) has introduced a scheduled hill climbing algorithm that leverages two distinct strategies: 1) Programmatic Neighborhood Generation, which involves generating a neighborhood of programs by selecting a node from the Abstract Syntax Tree (AST) of a given program and substituting it with a subtree that is randomly generated in accordance with the production rules and sampling strategies of a Domain-Specific Language (DSL); and 2) Latent Space Representation, where the neighborhood is defined in a latent space by training a variational autoencoder on a corpus of randomly generated DSL programs, thereby creating a more abstract and potentially more informative neighborhood structure. Additionally,  Jin et al. (2023) contribute to this line of research by defining neighborhood algorithms based on a minimum cosine similarity threshold of 60% between the embeddings of a query vector and code snippets, enhancing the search process by ensuring that algorithmic variations are semantically coherent and aligned with the underlying problem structure.
@@ 4.2.3. Gradient-based Search
Gradient-based search leverages a (pseudo) gradient direction in the generation of the new design in each iteration. Unlike searching in continuous space with an explicit differentiable objective formulation, calculating the gradient in the text and code space is challenging. As a result, an estimated descent direction is used as a pseudo gradient in the text and code space.
Pryzant et al. (2023) introduce automatic prompt optimization, which improves prompts automatically by adjusting them based on natural language “gradients” derived from training data. This approach has led to a substantial improvement in LLM performance on various NLP tasks and jailbreak detection by up to 31%. Moreover, Nie et al. (2024) explore the use of LLMs as interactive optimizers for solving maximization problems in a text space through natural language and numerical feedback. LLMs exhibit effectiveness when provided with directional feedback, which is an extension of first-order feedback in the natural language domain. Furthermore, Tang et al. (2024c) introduce a novel Gradient-inspired LLM-based Prompt Optimizer (GPO) that draws parallels with gradient-based model optimizers. By incorporating updated direction and method concepts from gradient-based optimization, GPO enhances prompt optimization strategies, leading to significant performance enhancements in experimental assessments. This integration of LLMs with gradient-based search methods showcases the potential for improving prompt optimization and overall LLM performance.
For an accurate gradient, Guo et al. (2024a) perform the gradient-based search in a mapped continuous space and proposes a collaborative optimization strategy that combines a gradient-based optimizer with an LLM for tackling complex non-convex optimization challenges. The LLM is utilized to derive solutions from natural language instructions, while the gradient-based optimizer performs locally optimal updates at each iteration. This collaborative approach leverages the strengths of both optimizers, with the inferred results from LLMs serving as starting points for subsequent stages of gradient optimization.

### v2

@@ 3.3. LLMs as Extractors (LLMaE)
LLMaE employs LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then utilized in the enhancement of algorithm-based problem solving (Fig 6). This process leverages the unique capabilities of LLMs, such as text and code comprehension, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods (Wu et al., 2024c).
@@
In the context of feature extraction, Kristiadi et al. (2024) explore the use of LLMs in accelerating Bayesian optimization in the molecular space. LLMs are used as fixed feature extractors for standard Bayesian optimization surrogate models, and parameter-efficient finetuning methods are leveraged to obtain the posterior of the LLM surrogate.
Beyond conventional feature extraction, LLMs excel in text comprehension and knowledge extraction. An example in LLM-assisted algorithm selection by Wu et al. (2024c) utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. In another innovative application, Du et al. (2024c) propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, Becker et al. (2024) discuss the use of LLMs in the automotive and supplier industries, specifically focusing on retrieval-augmented generation systems to improve information retrieval from technical documentation. Memduhoğlu et al. (2024) use LLM embeddings to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
The multi-modal capabilities of LLMs also distinguish them from traditional methods. Park et al. (2024) demonstrate that incorporating lexical information extracted from LLMs into an acoustic-based speaker diarization system through a probabilistic multi-modal decoding process and beam searches can significantly improve speech-processing performance.
@@ 3.4. LLMs as Designers (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 7). This utilization of LLMs extends their application beyond traditional boundaries, enabling them to actively participate in algorithm development by generating heuristics (Liu et al., 2024d), writing code snippets (Ma et al., 2024b), or formulating functions (Romera-Paredes et al., 2024) that can be integrated into algorithmic systems. By doing so, LLMs can significantly accelerate the algorithm design process, reduce human effort, and bring creativity and optimization to algorithm development (Liu et al., 2024d), which is difficult to achieve through conventional methods alone.
@@
Function design is among the early applications of LLMaD. Eureka (Ma et al., 2024b) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for RL. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. Similarly, Auto MC-Reward (Li et al., 2024b) utilizes LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. Moreover, FunSearch (Romera-Paredes et al., 2024) adopts LLMs for function generation in an evolutionary framework with a multi-iland population management. FunSearch demonstrated promising results on both mathematical problems and combinatorial optimization problems.
@@
EoH (Liu et al., 2024d), originates from AEL (Liu et al., 2023b), presents an early attempt to adopt LLM as a designer for automatic heuristic design. It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the algorithmic heuristic. EoH is applied to the key heuristic component design for combinatorial optimization problems, where it surpasses both traditional human-designed metaheuristics and deep-learning-based neural solvers. The application of EoH has expanded to include cost-aware acquisition function design for Bayesian optimization (Yao et al., 2024a), adversary attack algorithm design for image adversary attack (Guo et al., 2024b), and heuristic design for edge server task scheduling Yatong et al. (2024). LLaMEA (van Stein and Bäck, 2024) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (Ye et al., 2024a) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. With the reflection, better results on black-box cases (no problem-specific information is provided in the prompts) have been observed. Moreover, Yan et al. (2024) explore the use of LLMs to generate drafts of high-quality algorithms for electronic health records phenotyping, aiming to improve efficiency and scalability in observational health research and Hazra et al. (2024) combine LLMs with heuristic planning to generate cost-effective plans, showing superior performance in LLM planning approaches. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (Yao et al., 2024b) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
@@
LLM-based agent design has also gained much attention. For example, ADAS (Hu et al., 2024) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents. They present a novel algorithm, meta agent search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains. Further studies on LLM-based agent systems are discussed in (Sun et al., 2024a) and (Liu et al., 2024e).
@@
There has been a considerable effort on using LLMs for optimization modeling. Ahmed and Choudhury (2024) assess various LLMs for transforming linguistic descriptions into mathematical optimization problems, introducing a fine-tuning framework, LM4OPT, to enhance the performance of smaller models. Tang et al. (2024a) outline four key training dataset requirements for operational research LLMs and propose OR-Instruct, a semi-automated method for generating tailored synthetic data. Additionally, Mostajabdaveh et al. (2024) present a novel multi-agent framework that translates natural language into optimization models, utilizing relational identifier agents and a verification mechanism for improved assessment.
@@ 4. Search Methods
Incorporating LLMs within a search framework significantly enhances LLMs’ utility in algorithm design, making it the preferred option (Ma et al., 2024b; Liu et al., 2024d; Romera-Paredes et al., 2024). This section categorizes existing research according to the search methods employed, presents recent advancements, and discusses some of the limitations.
@@ 4.1. Sampling
The most straightforward search manner is by repeatedly instructing LLM to sample new designs (Zhou et al., 2022). The best sample is selected as the final design. However, recent studies have shown that simple sampling from LLMs can be costly (Zhang et al., 2024b).
@@ 4.1.1. Beam Search
Beam search explores multiple promising paths in the search space. For instance, in LLM-based prompt engineering (Zhou et al., 2022; Pryzant et al., 2023; Ma et al., 2024d), increasing the beam size significantly enhances performance (Zhou et al., 2022). Beyond prompt engineering, beam search is utilized in diverse applications: Text2Motion (Lin et al., 2023) employs it for robotic task planning, SayCanPay (Hazra et al., 2024) for action sequence planning in robots, enhancing speech-processing tasks, and in a confidence-aware method for structured data generation.
@@ 4.1.2. MCTS
Monte Carlo Tree Search (MCTS) is a tree search algorithm that uses random sampling to build a search tree and make decisions based on the results of these samples. It is particularly effective in problems with large and complex search spaces. For example, Dainese et al. (2024) propose GIF-MCTS, a code generation strategy using MCTS, in which nodes in the tree are programs and edges are actions. Each action taken from a parent node produces a new complete program. The upper confidence bound for trees is used to select which action to take. Similarly, VerMCTS (Brandfonbrener et al., 2024) constructs a search tree with progressive widening to effectively manage large action spaces defined by lines of code. In this search tree, expansion and evaluation are guided by the verifier, serving as a computationally inexpensive upper bound on the value function. Moreover, Wang et al. (2023b) regard prompt optimization as a strategic planning challenge, utilizing a principled planning algorithm based on MCTS to effectively explore the expert-level prompt space. The proposed PromptAgent generates precise insights and detailed instructions by analyzing model errors and providing constructive feedback.
@@ 4.2. Single-point-based Search
Single-point-based search methods iteratively refine a solution by leveraging neighborhood structures (Liu et al., 2024f) or specific search directions (Pryzant et al., 2023). While these methods can produce satisfactory results within a reasonable number of iterations, they may struggle with maintaining diversity and robustness in their search processes.
@@ 4.2.1. Hillclimb
Hillclimb search iteratively searches for better results, where the new one is generated based on the last one and the better one is survived. A basic version of Hillclimb is investigated by Zhang et al. (2024b) for algorithm design, where a heuristic is iteratively refined. The reasoning over the existing status is usually adopted to enhance each search step. For example, Li et al. (2024b) leverage LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards, which iteratively refines the reward function based on feedback from the agent’s interactions with the environment and Yang et al. (2023) develop a multi-module framework with innovative feedback mechanisms, demonstrating through both LLM-based and expert evaluations that LLMs can effectively produce scientific hypotheses that are both novel and reflective of reality.
@@ 4.2.2. Neighborhood Search
Neighborhood search methods explore the solution space by iteratively modifying a current solution to find an improved one in a structured neighborhood. Notably, the LLM-GS framework (Liu et al., 2024f) has introduced a scheduled neighborhood search method that leverages two distinct strategies: 1) Programmatic Neighborhood Generation, which involves generating a neighborhood of programs by selecting a node from the abstract syntax tree of a given program and substituting it with a subtree that is randomly generated following the production rules and sampling strategies of a domain-specific language; and 2) Latent Space Representation, where the neighborhood is defined in a latent space by training a variational autoencoder on a corpus of randomly generated DSL programs, thereby creating a more abstract and potentially more informative neighborhood structure. Additionally,  Jin et al. (2023) contribute to this line of research by defining neighborhood algorithms based on a minimum cosine similarity threshold of 60% between the embeddings of a query vector and code snippets, enhancing the search process by ensuring that algorithmic variations are semantically coherent and aligned with the underlying problem structure.
@@ 4.2.3. Gradient-based Search
Gradient-based search utilizes a (pseudo) gradient direction to generate new designs in each iteration. Unlike searches in a continuous space with a differentiable objective, calculating the gradient in the algorithm design space poses significant challenges. Consequently, a pseudo descent direction is often employed. Pryzant et al. (2023) introduce automatic prompt optimization, which improves prompts automatically by adjusting them based on natural language “gradients” derived from training data. Similarly, Tang et al. (2024c) introduce GPO, which draws parallels with gradient-based model optimizers for prompt optimization. Moreover, Nie et al. (2024) explore the use of LLMs as interactive optimizers for solving maximization problems in a text space through natural language and numerical feedback. For an accurate gradient, Guo et al. ([n. d.]) perform the gradient-based search in a mapped continuous space and proposes a collaborative optimization strategy that combines a gradient-based optimizer with an LLM for tackling complex non-convex optimization challenges. These works employ either pseudo gradients or gradients over mapped text space; however, a systematic study of gradients over general algorithm space is anticipated.